Find the number of units that must be produced and sold in order to yield the maximum profit, given the following equations for revenue and cost:



R(x) = 30x - 0.5x 2



C(x) = 5x + 7.




A. 32 units




B. 26 units




C. 35 units




D. 25 units

1
Expert's answer
2021-11-02T14:33:11-0400

"\displaystyle 
\text{The profit is given by }\\
p(x) = R(x)-C(x)\\
= 30x - 0.5x^2-5x-7\\
= -0.5x^2+25x-7\\
p'(x) = -x+25\\
\text{When profit is maximum, $p'(x) = 0$ }\\
\therefore x = 25"